OpenStudy (anonymous):

-6 + x/4 = -5 PLEASE HELP! THANKS! :)MEDAL

OpenStudy (anonymous):

-6 + x/4 = -5 x/4 = -5+6 x/4 = 1 x =???

OpenStudy (midhun.madhu1987):

can u complete the rest?